OpenStudy (anonymous):

2. 14x = 2 (5x + 14) (1 point)

OpenStudy (ajprincess):

First distribute the part in the right hand side. 14x=2(5x + 14) 14x=2*5x+2*14 14x=10x+28 Can u solve for x nw @HeyImKae?

OpenStudy (anonymous):

@ajprincess not really, I don't know how... D:

OpenStudy (ajprincess):

14x=10x+28 Plz follow the steps. !.Subtract 10x both sides. 2.Divide by 4 both sides.

OpenStudy (ajprincess):

can u do nw @HeyImKae?

OpenStudy (anonymous):

Is it 4.5? :3 @ajprincess

OpenStudy (ajprincess):

hmm no. 14x=10x+28 1.Subtract 10x both sides. 14x-10x=10x+28-10x 4x=28 2.Divide by 4 both sides. 4x/4=28/4 x=7 Is it clear?